# Traffic Engineering in Large-scale Networks with Generalizable Graph Neural Networks

### v1

@@ Abstract
Traffic engineering (TE) in large-scale computer networks has become a fundamental yet challenging problem, owing to the swift growth of global-scale cloud wide-area networks or backbone low-Earth-orbit satellite constellations. To address the scalability issue of traditional TE algorithms, learning-based approaches have been proposed, showing potential of significant efficiency improvement over state-of-the-art methods. Nevertheless, the intrinsic limitations of existing learning-based methods hinder their practical application: they are not generalizable across diverse topologies and network conditions, incur excessive training overhead, and do not respect link capacities by default.
This paper proposes TELGEN, a novel TE algorithm that learns to solve TE problems efficiently in large-scale networks, while achieving superior generalizability across diverse network conditions. TELGEN is based on the novel idea of transforming the problem of “predicting the optimal TE solution” into “predicting the optimal TE algorithm”, which enables TELGEN to learn and efficiently approximate the end-to-end solving process of classical optimal TE algorithms. The learned algorithm is agnostic to the exact network topology or traffic patterns, and can efficiently solve TE problems given arbitrary inputs and generalize well to unseen topologies and demands.
We trained and evaluated TELGEN on random and real-world networks with up to $5000$ nodes and $10^{6}$ links. TELGEN achieved less than $3\%$ optimality gap while ensuring feasibility in all cases, even when the test network had up to 20×20$\times$20 × more nodes than the largest in training. It also saved up to $84\%$ solving time than classical optimal solver, and could reduce training time per epoch and solving time by 22224444 orders of magnitude than latest learning algorithms on the largest networks.
Index Terms:
Traffic Engineering, Network Optimization, Learn-to-Optimize, Graph Neural Network, Machine Learning
@@ I Introduction
With the emergence of large-scale and dynamic networks, classical TE faces fundamental challenges in terms of scalability, responsiveness and performance. Wide-area networks (WANs), such as Cloud WAN or low-Earth-orbit (LEO) satellite backbone networks, typically consist of hundreds to thousands of nodes and tens of thousands to millions of links. These networks present prohibitively large TE problem sizes that existing optimization-based algorithms cannot catch up with. Some networks also exhibit frequent topology changes and dynamic demand variations, which requires repeated and frequent solving of the TE problem at small timescales [39]. State-of-the-art linear program (LP) solvers, widely used to solve TE problems, can lead to up to cubic complexity in the number of links and traffic flows, which is prohibitive for modern network sizes. Some decomposition-based algorithms have been proposed [2, 27], but can only utilize limited CPU parallelism for moderate speed-up with no asymptotic improvement, and usually have suboptimal performance [39].
Recently, machine learning (ML) has been explored to speed-up TE in large-scale networks. By using ML algorithms such as graph neural network (GNN) or reinforcement learning (RL), networks can automatically learn and adapt to traffic patterns, predict congestion, and optimize routing in real-time. This results in improved network performance, reduced latency, and better resource utilization [22][31][29][41], making ML a promising tool for modern TE.
However, despite early success, current learning-based TE algorithms have several inherit limitations, which significantly hinder their application in real-world networks. First, existing algorithms are trained on the same network topology as testing or inference, and can adapt to at best minor topology changes such as individual link or node failures [39, 30]. This leads to inferior generalizability of the trained models, making them perform poorly on unseen topologies (such as when deployed in a new network) or in face of major topology changes (such as regional failures). The most generalizable learning-based TE solution to-date can adapt to up to $20\%$ topology changes, such as local link failures or congestion. They also do not generalize to traffic patterns substantially different from the training data, such as during periods with abnormal traffic. If the network conditions change majorly after training completes, the models must be completely retrained, causing significant time and computational overhead.
Second, many algorithms require extensive training time and computation power. For instance, a state-of-the-art solution TEAL [39] needs to be trained on latest GPU for more than a week for a network with over $1700$ nodes, and another state-of-the-art solution HARP [3] encounters prohibitive memory usage when scaling to networks with over $1000$ nodes. This makes it very difficult to (re-)train models for different network scenarios and conditions. Furthermore, some models do not always respect TE constraints such as link capacities and may cause excessive violations in some situations, in which case computationally expensive classical algorithms, such as ADMM—the alternating direction method of multipliers, have to be used to ensure feasibility and avoid causing excessive congestion [39].
@@
To address these challenges, we propose TELGEN, a novel end-to-end learning-based TE algorithm for large-scale networks. TELGEN is designed to efficiently solve TE problems in networks with diverse sizes and shapes, traffic patterns and conditions without or with minimal re-training. To achieve both optimality and generalizability, TELGEN employs a flexible GNN architecture that is agnostic to the topology or traffic demands, and is trained with (step-by-step) supervised learning to mimic the most powerful classical TE algorithms with high efficiency and optimality. Specifically, by casting TE problems as LPs, TELGEN learns the detailed steps that a classical TE algorithm (such as the interior-point method or IPM) takes to solve smaller-scale TE problems till optimality, and then generalizes the learned algorithm to larger networks and problem sizes utilizing the invariance properties of GNNs. Employing strong supervision from the classical algorithm significantly improves convergence and reduces training overhead in terms of the amount of data and computation needed.
@@
We trained and evaluated TELGEN on a wide range of randomly generated and real-world network topologies. Experiments showed that TELGEN could achieve close-to-optimal TE (with $<\!3\%$ optimality gap while ensuring feasibility), was up to $7$ times faster than a state-of-the-art IPM solver, and achieved up to $2$ orders of magnitude reduction in training time per epoch and prediction time over TEAL and 4 orders of magnitude reduction over HARP on large network topologies. Further, TELGEN maintained negligible optimality gap even when the test networks were 222220×20$\times$20 × larger than the largest network in training, or when demand patterns varied largely between training and testing, demonstrating superior generalizability in face of diverse network scenarios and conditions.
To summarize, our main contributions are as follows:
@@
  • We propose a general supervised learning approach to efficiently learn close-to-optimal and generalizable GNN-based algorithms for solving large-scale TE problems, by mimicking classical TE algorithm on smaller instances during training, and testing on larger, diverse networks.
  • We conducted extensive simulations on randomly generated and real network topologies with up to $5000$ nodes. Experiments showed the superior optimality, training and inference efficiency, and generalizability of TELGEN w.r.t. state-of-the-art solver and learning-based solutions.
In the following, Sec. II discusses related works. Sec. III describes the system model and problem formulation. Sec. IV describes the proposed GNN architecture and the training algorithm for solving large-scale TE problems. Sec. V shows evaluation results. Sec. VI concludes this paper.
@@ II-A Traffic Engineering
The primary goal of TE is to direct traffic within a data network to meet traffic demands by optimizing a chosen performance metric [26]. State-of-the-art TE solutions rely on formulating it as an LP or a convex program, which can be solved by commercial solvers using standard optimization algorithms (such as IPM, simplex method or primal-dual methods). However, the swift expansion and addition of new data centers or the emergence of new networks such as large-scale LEO satellite networks have made the TE optimization decision making process exceedingly slow and impractical at larger scales. Nowadays, it takes commercial solvers several hours to allocate traffic on WANs comprising thousands of nodes [39]. As a result, WAN operators are urgently looking for ways to expedite TE optimization to match the rapid growth of their networks. Methods haven been proposed to parallelize LP solvers, which can only achieve a coarse-level parallelization over CPUs and at best a moderate speed-up and no asymptotical improvement [2, 27]. Approximation algorithms such as Fleischer’s algorithm [13] have also been a solution to speed up TE, which nevertheless compromises optimality and can lead to many iterations to converge.
@@ II-B Machine Learning for Traffic Engineering
Methods have emerged to utilize ML to address the scalability issue in TE, with some promising initial results on significantly accelerating TE optimization [35, 4, 24, 33]. By training on historical or randomly generated data, ML models can make accurate predictions that enhance the efficiency of network management. Early approaches began with using Convolutional Neural Networks (CNNs) to efficiently handle spatial dependencies in network data [25], focuses on predicting traffic matrices as inputs to traditional TE algorithms. Subsequent works have explored deep RL to solve various TE problems including TE automation, delay minimization, power-awareness and dynamic changes of traffic demands [22, 31, 29, 41]. While promising, simple (deep) RL algorithms are difficult to encode or learn detailed network information such as topology and capacity constraints, and can lead to unstable performance or frequent capacity violations in practice. Some works have used GNNs for network performance modeling or prediction which could assist TE [12, 11], but do not directly address the efficiency and scalability bottleneck of classical TE solvers.
In TEAL [39], the authors combined multi-agent RL (MARL) with a GNN-based feature extractor to learn network-specific information (such as topology and flows), and applied classical ADMM to ensure feasibility. The GNN-MARL model was trained with historical or randomly generated demands on the target network, and performed reasonably well on the same network with different but similar demands. However, TEAL can only be applied to the exact same set of source-destination (SD) pairs on the same network as training, and have to be re-trained from scratch even when a new node is added or an existing node is removed, significantly limiting its generalizability. It also suffers from significant training time and overhead due to the use of RL, and cannot satisfy link capacity constraints without substantial help from ADMM.
Compared to RL-based solutions, another approach is to directly learn a mapping from problem inputs (or historical demands) to the TE solution, such as DOTE [30] and FIGRET [23]. These algorithms directly learn and encode the network topology into the neural model parameters, and hence cannot generalize to topologies beyond the training network. Recently, a transferrable neural model called HARP was proposed for minimizing the maximum link utilization (MLU) in TE [3], which can adapt to topology variation with up to $20\%$ difference in links. HARP is specially designed to minimize MLU and cannot be adapted to other TE objectives without modifying the model architecture. The goal of HARP’s transferrable design is to consider continuous topology changes such as local link/node failures during TE operations, and it does not tackle the issue of generalizing predictions to networks significantly different from or larger than training. Compared to HARP, our proposed TELGEN can generalize to networks with up to 20×20$\times$20 × more nodes than those in training, owing to its principled design around generalizability.
@@ II-C Graph Neural Network
GNNs [37] have emerged as a powerful framework for machine learning on graph-structured data. Let a graph be represented as $G=(V,E)$, where $V$ is the node set and $E$ is the edge set of the graph. The neighbor node set of a node $v$ is defined as $N(v)=\{u\in V\,|\,(v,u)\in E\}$. Define the adjacency matrix $A^{n\times n}$ with $A_{ij}=1$ is $e_{(i,j)}\in E$, and 0 otherwise, with $n=|V|$. Let $h_{v}$ be the node attribute for $v\in V$. A GNN model consists of a fixed number of GNN layers, denoted by $K$, and let a node’s attribute after layer $k$ be $h_{v}^{k}$. In each layer $k$, a message passing function denoted as ${\sf msg}^{k}(\cdot)$ is applied to pass and aggregate the node attributes of all neighbors of each node $v$, and an aggregation function denoted as ${\sf aggr}^{k}(\cdot)$ is then applied to aggregate the aggregated neighbor messages with the node’s own attribute $h_{v}^{k-1}$, outputting the updated node attribute $h_{v}^{k}$. A GNN layer can thus be represented as:
$h_{v}^{k}={\sf aggr}^{k}(h_{v}^{k-1},{\sf msg}^{k}(\{h_{u}^{k-1}|u\in N(v)\})).$ (1)
To make a prediction, a readout function $r(\cdot)$ is further applied to node attributes after layer $K$ to predict the target node, link or graph-level labels. In a GNN, the functions ${\sf msg}^{k}$, ${\sf aggr}^{k}$ and $r$ are modeled by parameterized neural networks (such as multi-layer perceptrons or MLPs), whose parameters are trained with training data guided by a loss function.
@@ III System Model and Problem Formulation
The goal of TE is to allocate network resources among clients and optimize for network performance, such as maximizing throughput, minimizing latency, or minimizing maximum link utilization (MLU). The widely used TE formulation involves splitting traffic demands between source-destination (SD) pairs along multiple paths. This involves determining the best routes for data packets to travel and balancing the load across various network paths to avoid congestion and ensure efficient utilization of bandwidth. Without loss of generality, we assume the objective of the network operator is to maximize the overall network throughput when meeting the demands of network SD pairs, subject to the link capacities.
@@ III-B TE Optimization Problem Formulation
Objective. The basic goal of TE optimization is to maximize network flow for all demands. This lead to the objective function as: maximize $f_{0}(\mathcal{R})\triangleq\sum_{i\in\textit{D}}d_{i}\sum_{p\in P_{i}}\mathcal{R}(p)$. We use this throughput maximization objective as the primary example in our description and evaluation, although our designed solution should be adaptable to other linear objectives (such as minimizing MLU or routing cost), and possibly convex non-linear objectives as well (such as fairness).
Constraints. The problem has two sets of constraints. First, the total allocated traffic for each SD pair cannot exceed its demand, and hence $\sum_{p\in P_{i}}\mathcal{R}(p)\leq 1$ for every $i\in D$. Second, the traffic allocated on a link cannot exceed its capacity, and hence we have $\sum_{i\in D}d_{i}\sum_{p\in P_{i}:l\in p}\mathcal{R}(p)\leq C(l)$ for $\forall l\in L$.
@@
For simplicity, we define constraint functions $f^{D}_{i}(\mathcal{R})\triangleq\sum_{p\in P_{i}}\mathcal{R}(p)-1$, and $f^{L}_{l}(\mathcal{R})\triangleq\sum_{i\in D}d_{i}\sum_{p\in P_{i}:l\in p}\mathcal{R}(p)-C(l)$. We want to design an algorithm that solves (6) to its optimality, ensures feasibility of the solution, and can do so efficiently on very large network scenarios (with thousands of nodes) even on those not observed in the training process.
IV TELGEN: A Learning-based TE Framework

### v2

@@ Abstract
Traffic Engineering (TE) in large-scale networks like cloud Wide Area Networks (WANs) and Low Earth Orbit (LEO) satellite constellations is a critical challenge. Although learning-based approaches have been proposed to address the scalability of traditional TE algorithms, their practical application is often hindered by a lack of generalization, high training overhead, and a failure to respect link capacities. This paper proposes TELGEN, a novel TE algorithm that learns to solve TE problems efficiently in large-scale network scenarios, while achieving superior generalizability across diverse network conditions. TELGEN is based on the novel idea of transforming the problem of “predicting the optimal TE solution” into “predicting the optimal TE algorithm”, which enables TELGEN to learn and efficiently approximate the end-to-end solving process of classical optimal TE algorithms. The learned algorithm is agnostic to the exact underlying network topology or traffic patterns, and is able to very efficiently solve TE problems given arbitrary inputs and generalize well to unseen topologies and demands. We train and evaluate TELGEN with random and real-world topologies, with networks of up to $5000$ nodes and $3.6\times 10^{6}$ links in testing. TELGEN shows less than $3\%$ optimality gap while ensuring feasibility in all testing scenarios, even when the test network has $2\text{-}20\times$ more nodes than the largest training network. It also saves up to $84\%$ TE solving time than traditional interior-point method, and reduces up to $79.6\%$ training time per epoch than the state-of-the-art learning-based algorithm.
@@ I Introduction
Crucial challenges for the central controller to manage TE are diverse network conditions such as the traffic pattern and network topology changes, and scalability issues in large-scale networks. Depending on the network scenario, network traffic patterns are governed by time-varying user activities, automated machine tasks, natural, social and political events, and so on. The topology could also vary because of network congestion, outages, or unexpected failures, which can significantly affect network performance such as throughput and quality-of-service (QoS) that can further lead to higher operational costs, revenue loss or penalty, if not dealt with promptly. Tackling this challenge requires periodically re-configuring network TE in the time scale of minutes to hours, which poses a significant overhead on the central controller [43].
@@
The scale of modern computer networks exacerbates the challenge of TE. Current Wide-Area Networks (WANs), such as Cloud-WAN or Low Earth Orbit (LEO) satellite backbone network, typically consist of hundreds to thousands of nodes and tens of thousands to millions of links. These networks present prohibitively large TE problem sizes that existing optimization-based algorithms cannot catch up with. State-of-the-art Linear Program (LP) solvers, widely used to solve TE problems, can lead to up to cubic complexity in the number of links and traffic flows, which is prohibitive for modern network sizes.
Some decomposition-based algorithms have been proposed [1, 26], but can only utilize limited CPU parallelism to achieve up to moderate speed-up with no asymptotic improvement [43]. With the emergence of large-scale and dynamic networks[32, 47], classical TE faces fundamental challenges in terms of scalability and responsiveness. While traditional WAN TE cycles could tolerate latencies of several minutes, modern environments like large-scale data centers and Low-Earth Orbit satellite constellations now demand real-time responses on the order of seconds [44]. For instance, in LEO networks, the topology changes so rapidly that over 50% of network paths can become obsolete within 150 seconds [39]. State-of-the-art LP solvers, however, can require tens of seconds for computation alone, a timeframe that renders them ineffective for such dynamic scenarios [39]. This creates a critical need for a new class of TE solutions that can deliver near-optimal traffic allocations at millisecond latencies, thereby making real-time control loops feasible.
Recently, the power of machine learning (ML) has been explored to speed-up TE in large-scale networks. By using ML algorithms such as graph neural network (GNN) or reinforcement learning (RL), networks can automatically learn and adapt to traffic patterns, predict congestion, and optimize routing in real-time. This results in improved network performance, reduced latency, and better resource utilization [22][31][28][46], making ML a promising tool for modern TE. Despite the early success, however, the current learning-based TE algorithms have several inherit limitations, which significantly hinder their practical application in real-world TE scenarios. First, existing algorithms typically involve training with various traffic demands on the same network topology as testing or inference, with at best minor topology changes such as individual link or node failures [43, 30]. In [2], up to $20\%$ of the topology changes are allowed by the form of capacity variation considered as inactive links or subsets of nodes from one topology. This leads to inferior generalizability of the trained models, making them perform poorly on unseen topologies (such as when deployed in a new network) or in face of significant topology changes (such as regional failures); they also do not generalize to traffic matrices that are substantially different from the training data which is not flexible when dealing with training and testing data from different ranges. Second, many of these algorithms require extensive training time and computation power. For instance, TEAL [43] needs to be trained on the latest GPUs for more than a week for a large topology and HARP [2] has difficulty dealing with large topology with node number over $1000$ in the traffic matrix processing part. This makes it very difficult to (re-)train models for different network scenarios and conditions. Furthermore, as they adopt unsupervised or self‑supervised learning approaches, existing algorithms do not always respect TE constraints, for example link capacities, and must be combined with expensive classical algorithms such as Alternating Direction Method of Multipliers (ADMM) to ensure feasibility of the predicted TE solutions.
@@
To address the above challenges, we propose TELGEN, a novel end-to-end learning-based TE algorithm for large-scale networks. TELGEN is designed with the goal of enabling generalizability to diverse network topologies, sizes, traffic patterns and conditions without or with minimal re-training. To achieve both generalizability and constraint satisfaction, TELGEN employs a flexible GNN architecture that is agnostic to the topology or traffic demands, and is trained with (step-by-step) supervised learning to mimic the most powerful classical TE algorithms with high efficiency and optimality. Specifically, by casting TE problems as LPs, TELGEN learns the detailed steps that a classical TE algorithm (such as the interior-point method or IPM) takes to solve smaller-scale TE problems till optimality, and then generalizes the learned algorithm to larger networks and problem sizes utilizing the invariance properties of GNNs. By employing strong supervision from the classical algorithm, TELGEN reduces the training overhead in terms of the amount of data and computation needed and speeds-up training by several orders of magnitude.
@@
To evaluate its performance, we train and evaluate TELGEN on a wide range of randomly generated and real-world network topologies. Experiments show that TELGEN can achieve close-to-optimal TE (with $3\%$ optimality gap while ensuring feasibility), is up to $7$ times faster than a state-of-the-art IPM solver, and reduces the training time per epoch by up to $79.6\%$ over TEAL and HARP. Further, TELGEN maintains negligible optimality gap even when the test networks are up to $2$-$20\times$ larger than the largest network in training, or when demand patterns vary largely between training and testing, demonstrating superior generalizability of TELGEN in face of diverse network scenarios and conditions. Our main contributions are as follows:
@@
  • We propose a general supervised learning approach for efficiently learning close-to-optimal and generalizable GNN-based algorithms for solving large-scale TE problems, by mimicking classical TE algorithm on smaller instances for the training process and testing on larger, diverse networks.
  • We conduct extensive simulations on both randomly generated and real network topologies with up to $5000$ nodes, and demonstrate the superior optimality, training efficiency, inference time of TELGEN, compared to state-of-the-art solver and learning-based algorithms. Especially, the generalizability over network topology, size, condition as well as traffic demand of TELGEN has been justified, compared to state-of-the-art solver and learning-based algorithms.
In the following, Sec. II discusses related works. Sec. III describes the system model and problem formulation. Sec. IV describes the proposed GNN architecture and the training algorithm for solving large-scale TE problems. Sec. V shows evaluation results. Sec. VII concludes this paper.
@@ II-A Traffic Engineering
The primary goal of TE is to direct traffic within a data network to meet traffic demands by optimizing a chosen performance metric [25]. State-of-the-art WAN TE solutions rely on formulating it as LPs or convex programs, which can be solved by commercial solvers using standard optimization algorithms (such as IPM, simplex method or primal-dual methods). However, the swift expansion and addition of new data centers or the emergence of new networks such as large-scale LEO satellite networks have made the TE optimization decision making process exceedingly slow and impractical at larger scales. Nowadays, it takes commercial solvers several hours to allocate traffic on WANs comprising thousands of nodes [43]. As a result, WAN operators are urgently looking for ways to expedite TE optimization to match the rapid growth of their networks. Methods haven been proposed to parallelize LP solvers, which nevertheless only achieve a coarse-level parallelization over CPUs and at best a moderate speed-up [1, 26]. Approximation algorithms such as Fleischer’s algorithm [11] have also been a solution to speed up TE, which nevertheless compromises optimality and can lead to many iterations to converge.
@@ II-B Machine Learning for Traffic Engineering
Methods have emerged to utilize ML to address the scalability issue in TE, with some promising initial results on significantly accelerating TE optimization [37, 3, 24, 34]. By training on historical or randomly generated data, ML models can make accurate predictions that enhance the efficiency of network management. Early approaches began with using Convolutional Neural Networks (CNNs) to efficiently handle spatial dependencies in network data [24], focuses on predicting traffic matrices as inputs to traditional TE algorithms. Subsequent works have explored deep RL to solve various TE problems including TE automation, delay minimization, power-awareness and dynamic changes of traffic demands [22, 31, 28, 46]. While promising, simple (deep) RL algorithms are difficult to encode or learn detailed network information such as topology and capacity constraints, and can lead to unstable performance or frequent capacity violations in practice.
In TEAL [43], the authors combined multi-agent RL (MARL) with a GNN-based feature extractor to learn network-specific information (such as topology and flows), and applied classical ADMM to ensure feasibility. The GNN-MARL model is trained with historical or randomly generated demands on the target network, and performs reasonably well on the same network with different demands. However, TEAL can only be applied to the exact same set of source-destination (SD) pairs on the same network as training, and have to be re-trained from scratch even when a new node is added or an existing node is removed, significantly limiting its generalizability. It also suffers from significant training time and overhead due to the use of RL, and cannot satisfy link capacity constraints without substantial help from ADMM.
Compared to RL-based solution, another approach is to directly learn a mapping from the problem inputs (or historical demands) to the TE solution, such as in DOTE [30] and FIGRET [23]. These algorithms directly learn and encode the network topology into the neural model parameters, and hence cannot generalize well on topologies beyond the training set. A related recent paper presents a neural TE model, HARP, that can generalize to changing topologies [2] with a topology variation up to $20\%$ and the graph size up to less than 1000 nodes. HARP considers continuous topology changes during TE execution, and does not to tackle the issue of TE solution generalizability to topologies significantly different from or larger than those in the training dataset. As such, the approach would not be able to reduce training overhead by training on smaller networks and then applied on much larger ones, or instantly be applied when significant topology changes such as regional failures happen.
@@ II-C Graph Neural Network
GNNs [40] have emerged as a powerful framework for machine learning on graph-structured data. Let a graph be represented as $G=(V,E)$, where $V$ is the node set and $E$ is the edge set of the graph. The neighbor node set of a node $v$ is defined as $N(v)=\{u\in V|(v,u)\in E\}$. Define the adjacency matrix $A^{n\times n}$ with $A_{ij}=1$ is $e_{(i,j)}\in E$, and 0 otherwise, with $n=|V|$. Let $h_{v}$ be the node attribute for $v\in V$. A GNN model consists of a fixed number of GNN layers, denoted by $K$, and let a node’s attribute after layer $k$ be $h_{v}^{k}$. In each layer $k$, a message passing function denoted as $MSG^{k}(\cdot)$ is applied to pass and aggregate the node attributes of all neighbors of each node $v$, and an aggregation function denoted as $g^{k}(\cdot)$ is then applied to aggregate the aggregated neighbor messages with the node’s own attribute $h_{v}^{k-1}$, outputting the updated node attribute $h_{v}^{k}$. A GNN layer can thus be represented as:
$h_{v}^{k}=g^{k}(h_{v}^{k-1},MSG^{k}(\{h_{u}^{k-1}|u\in N(v)\})).$ (1)
To make a prediction, a readout function $r(\cdot)$ is further applied to node attributes after layer $K$ to predict the target node, link or graph-level labels. In a GNN, the functions $MSG^{k}$, $g^{k}$ and $r$ are modeled by parameterized neural networks (such as multi-layer perceptrons or MLPs), whose parameters are trained with training data guided by a loss function.
@@ III System Model and Problem Formulation
The goal of TE is to allocate network resources among clients and optimize for network performance, such as maximizing throughput, minimizing latency, or minimizing maximum link utilization (MLU). The widely used TE formulation involves splitting traffic demands between source-destination (SD) pairs along multiple paths. This involves determining the best routes for data packets to travel and balancing the load across various network paths to avoid congestion and ensure efficient utilization of bandwidth. Our work focuses on traffic engineering in large-scale point-to-point networks, which is motivated by challenges observed in practical settings such as data centers and wide-area communication systems. In general, we believe our methodology can be applied to wireless scenarios and will study this in our future work. Without loss of generality, we assume the objective of the network operator is to maximize the overall network throughput when meeting the demands of network SD pairs, subject to the link capacities.
@@ III-B TE Optimization Problem Formulation
Objective. The basic goal of TE optimization is to maximize network flow for all demands. This lead to the objective function as: maximize $f_{0}(\mathcal{R})\triangleq\sum_{i\in\textit{D}}d_{i}\sum_{p\in P_{i}}\mathcal{R}(p)$. We use this throughput maximization objective as the primary example in our description and evaluation, although our designed solution can be easily adopted to other linear objectives (such as minimizing MLU or routing cost), and possibly convex non-linear objectives as well (such as fairness).
Constraints. The problem has two sets of constraints. First, the total allocated traffic for each SD pair cannot exceed its demand, and hence $\sum_{p\in P_{i}}\mathcal{R}(p)\leq 1$ for every $i\in D$. Second, the traffic allocated on each link cannot exceed its capacity, and we have $\sum_{i\in D}d_{i}\sum_{p\in P_{i}:l\in p}\mathcal{R}(p)\leq C(l)$ for $\forall l\in L$.
@@
For simplicity, we define constraint functions $f^{D}_{i}(\mathcal{R})\triangleq\sum_{p\in P_{i}}\mathcal{R}(p)-1$, and $f^{L}_{l}(\mathcal{R})\triangleq\sum_{p\in P_{i}:l\in p}\mathcal{R}(p)-C(l)$. We are interested in designing an algorithm that solves (2) to its optimality, ensures feasibility of the solution, and can do so efficiently on very large network scenarios (with thousands of nodes) even on those not observed in the training process.
IV The TELGEN Learning-based TE Framework
Diverse network conditions such network topology changes and dynamic demands complicate the task of efficiently TE. Additionally, scalability issues arise as networks grow in size, making it difficult to apply traditional TE methods that struggle to adapt and perform in real-time. These challenges underscore the need for innovative solutions that can dynamically respond to fluctuating network conditions and scale seamlessly with network growth. Therefore, we design our TE solution with a focus on addressing these two critical challenges—diverse network conditions and scalability. Our goal is to develop a robust framework that enhances TE, ensuring efficient and reliable data flow in increasingly complex network environments.